Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR12 in R68 @ 15:30 pm on Tuesday 12th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • Actions from last meeting
    • DS readback current requires calibration (DL)
    • Install FC monitoring in ISIS MCR (Pierrick,Paolo)
    • Install SSH Enable/Disable mimic : Pierrick
  • DC report
  • Magnet & Channel commissioning
  • Operations in the next few days
    • Running plans
      • Wednesday (no beam so no running)
      • Thursday - run with FC only
        • morning - ramp FC
        • 10'ish (?) - 4pm : take data
        • 4 pm - ramp down FC
      • Friday - run with FC only
        • Afternoon shift training from about 2:30 - 5
      • Saturday/Sunday : straight track running
    • Shift training
      • Training on Friday
        • 10 : MOM talk and Safety talk by MOM (Paul Kyberd) and Andy Nichols
        • 11-1 : Control room familiarisation in the remote control room (Francois)
        • 2 : Meet the MOM at ISIS control room to : get radiation badge and get card access to R5.2
        • 2:30 - 5
          • Guide to the layout of the MLCR (Francois)
          • PPS Training (Controlled access) (Henry)
          • Hall walkthrough (Andy)
          • Beamline and Target training (Henry)
  • AOB
  • DONM : 19th July in CR16

Minutes

  • Present : SB, PS, PF, CR, PK, JB, MC, KR, CW, VB, CM, JC, MU
  • Actions
    • DS readback current requires calibration (DL) : Stands
    • Install FC monitoring in ISIS MCR (Pierrick,Paolo) : Stands
    • Install SSH Enable/Disable mimic : Pierrick : Stands
  • Magnet and channel commissioning
  • can reproduce quench protect trip operation with only one power supply on ECE
  • can reach final desired current and sit there for about an hour. Can see noise growing on trace for
  • combined voltage taps. Looks like a set of spikes with ringing over a period of 10s with spikes growning
  • in amplitude with time until the trip level is reached. Maybe coming from PS and its controller but don't
  • know yet. Could be many things.
  • PRY/Magnets are also moving by up to 4 mm and Josef wants to stop this. Josef wants to put spacers between
  • the SSs and the PRY. Jason is looking at this and has a design
  • but it won't be ready for this run. Crap - what do I do with shifters. Possibly could get this done for the end of
  • the next shutdown. **
  • Whats the plan? We run the ECE with a single PS tomorrow and look at ringing behaviour. Whilst ringing adjust
  • PS behaviour to see what effect this has.
  • Victoria : is there any historical data from when SSU/D were powered together as there was less movement then?
  • Mike : noone's looked.
  • JC : can see a 0.1 Hz signal on FC all the time from somewhere. When we see a spike on SSU we also see a spike on the
  • FC.
  • Trips are close together - growth of spikes. Time will be needed.
  • Plots look OK
  • JC : can the tracker see if the SSs have shifted.
  • MU : Paul was going to look at it in simulation.
  • Paul : haven't got too far. Compare tracker distributions before and after powering and see if there is any difference.
  • JC wrote a document about running the FC and has a document for shifters.

Updated by Boyd, Steven almost 6 years ago ยท 7 revisions